QUESTION 15 Compute ?*AP given 4+ [* :]md P=[^;]
Answer #1

P5 [] Det p= 5.2 - (2760) = 10-8 PART 14 J[ 2 ] 11 -16 +12 10+12 755 -27 1 -32+30 20+30 ) \: -4.2] - 22755-27 21 -2 50 1-4 2]